OpenStudy (anonymous):

solve the inequality and specify the answer using interval notation 9/10<3x-1/-2<91/100

OpenStudy (anonymous):

\[\frac{9}{10}<\frac{3x-1}{-2}<\frac{91}{100}\] Is this your question?

OpenStudy (anonymous):

yes

OpenStudy (rogue):

You must solve each side separately.\[.9 < \frac {3x - 1}{-2}\]\[\frac {3x - 1}{-2} < .91\]Then combine the interval solution for x.

OpenStudy (anonymous):

Multiply each side by -2, reversing the inequality signs: \[\frac{-18}{10}>3x-1>\frac{-182}{100}\] Rewrite it with less than signs: \[\frac{-91}{50}<3x-1<\frac{-9}{5}\] Add 1 to both sides: \[\frac{-41}{50}<3x<\frac{-4}{5}\] Divide each side by 3: \[\frac{-41}{150}<x<\frac{-4}{15}\] Interval notation: \[(-\frac{41}{50},-\frac{4}{15})\]

OpenStudy (anonymous):

thank you